# FootFormer: Estimating Stability from Visual Input

### v1

@@ 1 Introduction
Despite extensive work on estimating human body pose and motion [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12], significantly less attention has been paid to inferring physical quantities such as foot pressure and stability. Consider the Center of Pressure (CoP), which marks the net point of reactive force between a person and the ground plane. CoP’s location relative to the whole body center of mass has been identified as a determinant of stability in human motion [13, 14, 15]. Typically, CoP is obtained in a lab setting by force plates or insole foot pressure sensors [16, 17]. Yet, as CoP is correlated with whole body kinematics, specifically mass and acceleration of body parts [18], estimating it visually in natural settings is plausible [19, 20]. Vision models that infer motion dynamics quantities like CoP may enable scalable video-based analyses of human balance and stability, with applications in kinesiology, animation, and biomedical analysis. While prior work has explored video-based ground contact force estimation [17, 16, 21], the methods often only regress global scalar forces, omitting the rich structure of foot-ground interaction. Recently, large datasets such as PSU-TMM100 [19] and MMVP [22] have collected synchronized video, motion capture (MoCap), and high-resolution insole foot pressure data. However, research on these datasets is often limited to regressing a single modality from single-frame input [19, 20] or to augment 3D pose estimation methods [22].
@@
    We validate FootFormer on PSU-TMM100 [19], MMVP [22], UnderPressure [17], and a newly collected Ordinary Movements dataset, and demonstrate FootFormer’s efficacy compared to other methods in achieving significantly better or equivalent performance on all three output forms, especially its statistically significant SOTA performance on stability estimation (Table 4).
@@
PNS [19] $\times$ $\times$
FPP-Net [22] $\times$
UP [17] $\times$
CoMNet [20] $\times$ $\times$ $\times$
@@ 2.1 From Kinematics to Ground Contact Dynamics
Prior works have explored estimating contact forces from kinematic and video inputs [23, 24, 25, 26, 27, 21], typically estimating simple vertical ground reaction forces (vGRFs) or binary foot contact, unlike the dense pressure distributions or foot-region contact used in our method. Dynamics constraints are often applied in postprocessing [28, 29, 30] to enforce physically plausible solutions. Alternatively, [31, 32] interleave kinematic predictions with physics-based simulation on a causal, frame-by-frame basis, designing and learning humanoid controllers in simulation [33, 31]. Other studies [34, 17, 16] analyze dynamics by observing MoCap data to estimate motion dynamics and exterior forces. While similar, our objective is to enable stability estimation with a more complete approximation of motion-stability that includes foot pressure, yielding center of pressure, base of support, foot contact, and center of mass (Table 1). More recent work [35] utilizes an LSTM to predict a scalar gait stance interpolation value for exoskeleton control. Similarly, [22] proposes a GRU-based network to estimate foot pressure and contact to augment 3D pose estimation. Conversely, in [36], a transformer is used to predict hip and knee joint angles given plantar pressure for purposes of exoskeleton control. In a similar vein, we utilize spatiotemporal pose inputs but focus attention on estimating dynamics that determine human stability.
@@ 2.2 Measuring Human Stability and Balance
Humans naturally sense and maintain balance [37], and the human visual cortex is attuned to observing other people’s balance and stability [38]. Quantitative evaluation of stability in research and clinical applications often use force plates to capture 3D forces for each foot while capturing body movements with MoCap technology [39, 40, 41]. A broad selection of mathematical models have been developed for stability, and a wide set of stability metrics have been defined in the literature [42, 14]. Recently reported works, including novel pendulum models [43] and deep learning for "On-Demand Balance Evaluation" [44, 45], are almost all limited to gait movements, synthesis (simulation)-oriented, dependent on lab force plates and MoCap systems, and most important, do not take video as a primary input.
Scott et al. [19] proposed PressNet-Simple to estimate foot pressure distributions and subsequently compute Center of Pressure (CoP) and Base of Support (BoS) on PSU-TMM100 dataset, which contains a large variety of pose orientations, two key components for stability analysis. Later work [20] added estimation of 3D body Center of Mass (CoM) to compute two classic stability measures, CoM-CoP and CoM-BoS. Du et al. [46, 47] use predominantly frontal pose sequences with both feet stationary on the ground to estimate CoP measures such as path length and sway area, which can indicate balance problems.
@@ 3.1 PSU-TMM100
PSU-TMM100 is a multimodal dataset of 100 human motion sequences (each 5min long) in which 10 participants perform 24-form Taiji (Tai Chi) (Figure 2 top 2 rows) [19]. PSU-TMM100 includes spatiotemporally-synchronized measurements of foot pressure insoles [48], MoCap markers [49], and two RGB video views. Because we are primarily interested in estimating foot pressure directly from vision, we use predicted OpenPose [1] 2D and 3D-triangulated joint positions provided in the dataset, instead of the MoCap data.
We follow the preprocessing proposed with PSU-TMM100: the OpenPose 2D and 3D joints are centered about the hip and z-score normalized per joint dimension. The raw pressure data is first clipped to the insole’s recording range (0-862 kPa). Then, because we predict foot pressure distributions as opposed to absolute pressure, we divide each frame by the total pressure. This removes the need for our model to implicitly learn to estimate subject weight and instead focus on spatial distribution of pressure. To define foot contact regions, we divide contact maps into $N$ uniform regions and classify a region in contact if the maximum pressure value exceeds a given threshold (10kPa in our case). Ground truth CoM is derived from Vicon’s Plug-in-Gait model [49].
@@ 3.3 UnderPressure
UnderPressure (UP) is a MoCap-foot contact synchronized dataset [17]. It consists of 10 participants performing a range of activities including locomotion, sitting, and interacting with objects such as stairs. The dataset uses an Xsens MVN MoCap system [50], providing 3D locations of 23 joints, and Moticon OpenGo Sensor Insoles [51] consisting of 16 plantar pressure sensors. UnderPressure evaluates pressure detection on both region contact and vertical ground reaction force (vGRF).
We use the preprocessing steps from UnderPressure [17]: joint positions are computed through forward kinematics from joint angles, with data augmentation applied via skeleton morphology variations using precomputed SVD basis vectors. Binary contact labels for the heel and toe-region of each foot are determined by whether the sum of smoothed vGRFs within those regions exceed a threshold of $5\%$ for the respective subject’s body weight.
@@ 3.4 MMVP
The dataset closest to PSU-TMM100 is the Multimodal MoCap Dataset with Vision and Pressure (MMVP) [22]. MMVP is composed of 16 participants and provides synchronized Azure Kinect [52] RGBD video and Xsensor pressure insole [53] data. MMVP provides mostly short (10 second maximum) motions, including dancing, jumping, and other exercises. The dataset provides relatively dense insole pressure maps ($\sim 500$ pressure sensors), foot contact labels extracted from 3D body meshes, accurate 2D and 3D body representations of dynamics, and comparatively fast-paced motions to that of PSU-TMM100.
We follow the preprocessing steps originally proposed in MMVP [22] for obtaining foot contact maps. For each frame, the insole pressure data is normalized to [0,1] by first dividing the pressure data by the respective subject’s body weight and applying a Sigmoid operation to the weight-normalized pressure. Foot contact labels are then determined (empirically [22]) by setting the contact threshold to 0.5. For 2D keypoint extraction, we opt to use OpenPose [1] instead of RTMPose [54].
@@ 4.1 Problem Formulation
Motion can be represented as a temporal sequence $S=\{x_{i}\}_{i=1}^{T}$ where $x_{i}$ denotes a pose, at time step $i$, represented as 2D joint coordinates extracted using OpenPose [1] keypoints, and $T$ is the sequence length. Given a sequence $S$ centered on the target pose $x_{t}$, FootFormer regresses three stability-related modalities: foot pressure distributions $P_{t}\in\mathbb{R}^{P^{\prime}}$, foot contact maps $C_{t}\in\{0,1\}^{N}$, and the 3D center of mass $m_{t}\in\mathbb{R}^{3}$. The pressure map $P_{t}$ contains $P^{\prime}$ flattened pressure values across both feet, while the contact map $C_{t}$ indicates binary contact states for $N$ discrete regions across both feet. We empirically set $T=9$, allowing the model to leverage four frames before and after the target frame.
@@ 4.2 Architecture
Graph Convolutional Network (GCN): We encode spatial structures using a GCN with learnable connectivity [55]. Each pose $x_{i}$ in the input sequence $S=\{x_{i}\}_{i=1}^{T}$ is represented as a fully connected graph with $K$ joints and weighted adjacency matrix $A\in\mathbb{R}^{K\times K}$. For input features $X_{in}\in\mathbb{R}^{K\times F}$ where $F$ is the joint feature dimension, the GCN outputs $X_{out}=AX_{in}W$ using learnable weights $W\in\mathbb{R}^{F\times d}$. Applied to each frame, this produces spatially-enhanced embeddings $E\in\mathbb{R}^{T\times d}$, where $d$ is the embedding dimension (512). To encode temporal structure, we add positional encodings $E=E+P$.
@@ 5.1 Implementation Details
Training Protocol: For PSU-TMM100 evaluation, we follow Scott et al.’s Leave-One-Subject-Out (LOSO) cross-validation scheme [19], training FootFormer on 9 subjects and testing on the remaining, left out subject, 10 times in round robin fashion. For the UnderPressure [17] and MMVP [22] datasets’ foot contact evaluation, we train our model from scratch using their respective original training protocols and code implementations.
Model Training: All models are trained on an Nvidia A6000 GPU with a batch size of 512. For FootFormer, we optimize our multi-task loss ($\lambda_{p},\lambda_{c},\lambda_{com}=1$) using AdamW, while baselines use Adam with $\mathcal{L}_{p}$ for pressure prediction and $\mathcal{L}_{c}$ for contact (FPP-Net [22] only), following their original protocols. Model-specific hyperparameters are in the Supplementary Material.
@@ 5.2 Evaluation
Baselines: We compare FootFormer trained on PSU-TMM100 to the following prior works: (1) PNS [19], a 4-layer fully connected network with added residual connections; (2) UP [17] (dubbed after its dataset), a 1D CNN + MLP model originally used to estimate vGRFs from pose sequences; and (3) FPP-Net [22], which encodes pose via a 1D CNN encoder before using a GRU to handle the sequential data followed by a dual-headed MLP regressor which jointly predicts foot pressure and binary contact. We adapt baselines minimally to ensure a fair comparison, only adjusting input and output sizes to fit PSU-TMM100. Table 1 provides an overview of model output capabilities.
Metrics: We report performance for three modalities: foot pressure, binary foot contact, and 3D center of mass (CoM). For foot pressure, we are interested in quantifying the normalized pressure distribution that facilitates Center of Pressure estimation, we report KLD distance of the predicted and ground-truth pressure distribution. We follow standard foot contact evaluation practice to report precision, recall, F1 score, and Intersection over Union (IoU) between the ground truth and predicted contact points. When evaluating Center of Mass, we use Euclidean error between our predicted CoM points and 3D CoM points provided in PSU-TMM100 measured with a Vicon MoCap system. FootFormer is the only model which jointly optimizes and outputs all three modalities (Table 1), directly enabling the quantification of human stability metrics [56, 57, 42].
@@ 5.3 Foot Pressure
Method PSU-TMM100 KLD $\downarrow$ OM KLD $\downarrow$
2D 3D 2D 3D
PNS [19] 2.82 $\pm$ 0.86 2.68 $\pm$ 0.94 3.53 $\pm$ 1.23 2.59 $\pm$ 0.94
FPP-Net [22] 1.40 $\pm$ 0.32 1.60 $\pm$ 0.48 1.52 $\pm$ 0.37 1.54 $\pm$ 0.34
UP [17] 1.45 $\pm$ 0.35 1.50 $\pm$ 0.35 1.69 $\pm$ 0.34 1.48 $\pm$ 0.41
@@
Table 2: Foot pressure estimation evaluated using KL Divergence (KLD) on PSU-TMM100 [19] and Ordinary Movements (OM) datasets. Results are averaged across all subjects using leave-one-subject-out (LOSO) cross-validation. Bold indicates the best (lowest KLD); denotes a statistically significant difference from FootFormer (Ours) under paired t-test (p < 0.05). FootFormer performs statistically significantly better or equivalently across the two datasets.

### v2

@@ 1 Introduction
Despite extensive work on estimating human body pose and motion [openpose, ChenNIPS14, chu2017multi, iqbal2017posetrack, JainTLB14, papandreou2017towards, PfisterCZ15, song2017thin, DeepPose14, convposemachines2016, Newell2016, Bulat2016], significantly less attention has been paid to inferring physical quantities such as foot pressure and stability. Consider the Center of Pressure (CoP), which marks the net point of reactive force between a person and the ground plane. CoP’s location relative to the whole body center of mass has been identified as a determinant of stability in human motion [Hof2007, Hof2008, Pai2003]. Typically, CoP is obtained in a lab setting by force plates or insole foot pressure sensors [GroundLink, UnderPressure]. Yet, as CoP is correlated with whole body kinematics, specifically mass and acceleration of body parts [winter2009biomechanics], estimating it visually in natural settings is plausible [ScottECCV2020, JesseRehabJournal]. Vision models that infer motion dynamics quantities like CoP may enable scalable video-based analyses of human balance and stability, with applications in kinesiology, animation, and biomedical analysis. While prior work has explored video-based ground contact force estimation [UnderPressure, GroundLink, li2019motionforcesfromvideo], the methods often only regress global scalar forces, omitting the rich structure of foot-ground interaction. Recently, large datasets such as PSU-TMM100 [ScottECCV2020] and MMVP [mmvp] have collected synchronized video, motion capture (MoCap), and high-resolution insole foot pressure data. However, research on these datasets is often limited to regressing a single modality from single-frame input [ScottECCV2020, JesseRehabJournal] or to augment 3D pose estimation methods [mmvp].
@@
    We validate FootFormer on PSU-TMM100 [ScottECCV2020], MMVP [mmvp], UnderPressure [UnderPressure], and a newly collected Ordinary Movements dataset, and demonstrate FootFormer’s efficacy compared to other methods in achieving significantly better or equivalent performance on all three output forms, especially its statistically significant SOTA performance on stability estimation (Table 4).
@@
PNS [ScottECCV2020] $\times$ $\times$
FPP-Net [mmvp] $\times$
UP [UnderPressure] $\times$
CoMNet [JesseRehabJournal] $\times$ $\times$ $\times$
@@ 2.1 From Kinematics to Ground Contact Dynamics
Prior works have explored estimating contact forces from kinematic and video inputs [brubakerContactDynamics, vondrakPhysicalSimulation, brubakerAnthropomorphicWalker, VideoMocapWeiChai2010, DataDrivenInverseDynamicsSiggraphAsia, li2019motionforcesfromvideo], typically estimating simple vertical ground reaction forces (vGRFs) or binary foot contact, unlike the dense pressure distributions or foot-region contact used in our method. Dynamics constraints are often applied in postprocessing [RempeContactDynamics2020, PhysCap2020, ShimadaPhysionical] to enforce physically plausible solutions. Alternatively, [SimPOEcvpr21, dynamicsRegulatedEgopose] interleave kinematic predictions with physics-based simulation on a causal, frame-by-frame basis, designing and learning humanoid controllers in simulation [VondrakMocapBipedControl2012, SimPOEcvpr21]. Other studies [Zell2020HumDynamics, UnderPressure, GroundLink] analyze dynamics by observing MoCap data to estimate motion dynamics and exterior forces. While similar, our objective is to enable stability estimation with a more complete approximation of motion-stability that includes foot pressure, yielding center of pressure, base of support, foot contact, and center of mass (Table 1). More recent work [Lhoste2024DeepLearningEO] utilizes an LSTM to predict a scalar gait stance interpolation value for exoskeleton control. Similarly, [mmvp] proposes a GRU-based network to estimate foot pressure and contact to augment 3D pose estimation. Conversely, in [RenExoskeletonTransformer], a transformer is used to predict hip and knee joint angles given plantar pressure for purposes of exoskeleton control. In a similar vein, we utilize spatiotemporal pose inputs but focus attention on estimating dynamics that determine human stability.
@@ 2.2 Measuring Human Stability and Balance
Humans naturally sense and maintain balance [Asslander_2015], and the human visual cortex is attuned to observing other people’s balance and stability [Firestone_2016]. Quantitative evaluation of stability in research and clinical applications often use force plates to capture 3D forces for each foot while capturing body movements with MoCap technology [Chaudhry_2011, chaudhry_2004, POPOVIC_2000]. A broad selection of mathematical models have been developed for stability, and a wide set of stability metrics have been defined in the literature [Bruijn_2013, Hof2008]. Recently reported works, including novel pendulum models [JessicaInvertedPendulum] and deep learning for "On-Demand Balance Evaluation" [wei2020using, mihalec2022balance], are almost all limited to gait movements, synthesis (simulation)-oriented, dependent on lab force plates and MoCap systems, and most important, do not take video as a primary input.
Scott et al. [ScottECCV2020] proposed PressNet-Simple to estimate foot pressure distributions and subsequently compute Center of Pressure (CoP) and Base of Support (BoS) on PSU-TMM100 dataset, which contains a large variety of pose orientations, two key components for stability analysis. Later work [JesseRehabJournal] added estimation of 3D body Center of Mass (CoM) to compute two classic stability measures, CoM-CoP and CoM-BoS. Du et al. [MultiTaskCoPGraphConvNetworks, ViewInvariantCoP] use predominantly frontal pose sequences with both feet stationary on the ground to estimate CoP measures such as path length and sway area, which can indicate balance problems.
@@ 3.1 PSU-TMM100
PSU-TMM100 is a multimodal dataset of 100 human motion sequences (each 5min long) in which 10 participants perform 24-form Taiji (Tai Chi) (Figure 2 top 2 rows) [ScottECCV2020]. PSU-TMM100 includes spatiotemporally-synchronized measurements of foot pressure insoles [tekscan_2020], MoCap markers [Vicon_2020], and two RGB video views. Because we are primarily interested in estimating foot pressure directly from vision, we use predicted OpenPose [openpose] 2D and 3D-triangulated joint positions provided in the dataset, instead of the MoCap data.
We follow the preprocessing proposed with PSU-TMM100: the OpenPose 2D and 3D joints are centered about the hip and z-score normalized per joint dimension. The raw pressure data is first clipped to the insole’s recording range (0-862 kPa). Then, because we predict foot pressure distributions as opposed to absolute pressure, we divide each frame by the total pressure. This removes the need for our model to implicitly learn to estimate subject weight and instead focus on spatial distribution of pressure. To define foot contact regions, we divide contact maps into $N$ uniform regions and classify a region in contact if the maximum pressure value exceeds a given threshold (10kPa in our case). Ground truth CoM is derived from Vicon’s Plug-in-Gait model [Vicon_2020].
@@ 3.3 UnderPressure
UnderPressure (UP) is a MoCap-foot contact synchronized dataset [UnderPressure]. It consists of 10 participants performing a range of activities including locomotion, sitting, and interacting with objects such as stairs. The dataset uses an Xsens MVN MoCap system [schepers2018xsens], providing 3D locations of 23 joints, and Moticon OpenGo Sensor Insoles [moticon2021opengo] consisting of 16 plantar pressure sensors. UnderPressure evaluates pressure detection on both region contact and vertical ground reaction force (vGRF).
We use the preprocessing steps from UnderPressure [UnderPressure]: joint positions are computed through forward kinematics from joint angles, with data augmentation applied via skeleton morphology variations using precomputed SVD basis vectors. Binary contact labels for the heel and toe-region of each foot are determined by whether the sum of smoothed vGRFs within those regions exceed a threshold of $5\%$ for the respective subject’s body weight.
@@ 3.4 MMVP
The dataset closest to PSU-TMM100 is the Multimodal MoCap Dataset with Vision and Pressure (MMVP) [mmvp]. MMVP is composed of 16 participants and provides synchronized Azure Kinect [microsoft2025azurekinect] RGBD video and Xsensor pressure insole [xsensor2025gaitmotion] data. MMVP provides mostly short (10 second maximum) motions, including dancing, jumping, and other exercises. The dataset provides relatively dense insole pressure maps ($\sim 500$ pressure sensors), foot contact labels extracted from 3D body meshes, accurate 2D and 3D body representations of dynamics, and comparatively fast-paced motions to that of PSU-TMM100.
We follow the preprocessing steps originally proposed in MMVP [mmvp] for obtaining foot contact maps. For each frame, the insole pressure data is normalized to [0,1] by first dividing the pressure data by the respective subject’s body weight and applying a Sigmoid operation to the weight-normalized pressure. Foot contact labels are then determined (empirically [mmvp]) by setting the contact threshold to 0.5. For 2D keypoint extraction, we opt to use OpenPose [openpose] instead of RTMPose [jiang2023rtmpose].
@@ 4.1 Problem Formulation
Motion can be represented as a temporal sequence $S=\{x_{i}\}_{i=1}^{T}$ where $x_{i}$ denotes a pose, at time step $i$, represented as 2D joint coordinates extracted using OpenPose [openpose] keypoints, and $T$ is the sequence length. Given a sequence $S$ centered on the target pose $x_{t}$, FootFormer regresses three stability-related modalities: foot pressure distributions $P_{t}\in\mathbb{R}^{P^{\prime}}$, foot contact maps $C_{t}\in\{0,1\}^{N}$, and the 3D center of mass $m_{t}\in\mathbb{R}^{3}$. The pressure map $P_{t}$ contains $P^{\prime}$ flattened pressure values across both feet, while the contact map $C_{t}$ indicates binary contact states for $N$ discrete regions across both feet. We empirically set $T=9$, allowing the model to leverage four frames before and after the target frame.
@@ 4.2 Architecture
Graph Convolutional Network (GCN): We encode spatial structures using a GCN with learnable connectivity [mao2019learning]. Each pose $x_{i}$ in the input sequence $S=\{x_{i}\}_{i=1}^{T}$ is represented as a fully connected graph with $K$ joints and weighted adjacency matrix $A\in\mathbb{R}^{K\times K}$. For input features $X_{in}\in\mathbb{R}^{K\times F}$ where $F$ is the joint feature dimension, the GCN outputs $X_{out}=AX_{in}W$ using learnable weights $W\in\mathbb{R}^{F\times d}$. Applied to each frame, this produces spatially-enhanced embeddings $E\in\mathbb{R}^{T\times d}$, where $d$ is the embedding dimension (512). To encode temporal structure, we add positional encodings $E=E+P$.
@@ 5.1 Implementation Details
Training Protocol: For PSU-TMM100 evaluation, we follow Scott et al.’s Leave-One-Subject-Out (LOSO) cross-validation scheme [ScottECCV2020], training FootFormer on 9 subjects and testing on the remaining, left out subject, 10 times in round robin fashion. For the UnderPressure [UnderPressure] and MMVP [mmvp] datasets’ foot contact evaluation, we train our model from scratch using their respective original training protocols and code implementations.
Model Training: All models are trained on an Nvidia A6000 GPU with a batch size of 512. For FootFormer, we optimize our multi-task loss ($\lambda_{p},\lambda_{c},\lambda_{com}=1$) using AdamW, while baselines use Adam with $\mathcal{L}_{p}$ for pressure prediction and $\mathcal{L}_{c}$ for contact (FPP-Net [mmvp] only), following their original protocols. Model-specific hyperparameters are in the Supplementary Material.
@@ 5.2 Evaluation
Baselines: We compare FootFormer trained on PSU-TMM100 to the following prior works: (1) PNS [ScottECCV2020], a 4-layer fully connected network with added residual connections; (2) UP [UnderPressure] (dubbed after its dataset), a 1D CNN + MLP model originally used to estimate vGRFs from pose sequences; and (3) FPP-Net [mmvp], which encodes pose via a 1D CNN encoder before using a GRU to handle the sequential data followed by a dual-headed MLP regressor which jointly predicts foot pressure and binary contact. We adapt baselines minimally to ensure a fair comparison, only adjusting input and output sizes to fit PSU-TMM100. Table 1 provides an overview of model output capabilities.
Metrics: We report performance for three modalities: foot pressure, binary foot contact, and 3D center of mass (CoM). For foot pressure, we are interested in quantifying the normalized pressure distribution that facilitates Center of Pressure estimation, we report KLD distance of the predicted and ground-truth pressure distribution. We follow standard foot contact evaluation practice to report precision, recall, F1 score, and Intersection over Union (IoU) between the ground truth and predicted contact points. When evaluating Center of Mass, we use Euclidean error between our predicted CoM points and 3D CoM points provided in PSU-TMM100 measured with a Vicon MoCap system. FootFormer is the only model which jointly optimizes and outputs all three modalities (Table 1), directly enabling the quantification of human stability metrics [jian1993trajectory, lugade2011center, Bruijn_2013].
@@ 5.3 Foot Pressure
Dataset (Metric) PSU-TMM100 (KLD $\downarrow$) OM (KLD $\downarrow$)
Method 2D 3D 2D 3D
PNS [ScottECCV2020] 2.82 $\pm$ 0.86 2.68 $\pm$ 0.94 3.53 $\pm$ 1.23 2.59 $\pm$ 0.94
FPP-Net [mmvp] 1.40 $\pm$ 0.32 1.60 $\pm$ 0.48 1.52 $\pm$ 0.37 1.54 $\pm$ 0.34
UP [UnderPressure] 1.45 $\pm$ 0.35 1.50 $\pm$ 0.35 1.69 $\pm$ 0.34 1.48 $\pm$ 0.41
@@
Table 2: Foot pressure estimation evaluated using KL Divergence (KLD) on PSU-TMM100 [ScottECCV2020] and Ordinary Movements (OM) datasets. Results are averaged across all subjects using leave-one-subject-out (LOSO) cross-validation. Bold indicates the best (lowest KLD); denotes a statistically significant difference from FootFormer (Ours) under paired t-test (p < 0.05). FootFormer performs statistically significantly better or equivalently across the two datasets.